OpenStudy (anonymous):

What are the roots of 2x+6=−5/x? Multiple choice A.3−+i2 B.−3−+i2 C.−3−+2i4 D.3−+i2√4

OpenStudy (anonymous):

@ehuman

OpenStudy (superdavesuper):

Multiply both sides by x u get 2x^2 + 6x = -5 or 2x^2 +6x +5 =0 You can then use the quadratic equation solution to solve this.

OpenStudy (anonymous):

I dont know how

OpenStudy (superdavesuper):

Have u see this formula before: ax^2 + bx + c = 0? Go check ur book if u are not sure.

OpenStudy (anonymous):

yes

OpenStudy (superdavesuper):

OK the u should know the solutions for that equation are (-b +/- SQRT(b^2-4ac))/2a

OpenStudy (anonymous):

Are you sure that your original equation is correct? I get something different than your answers